A ballistic pendulum malfunctions, and the bullet collides elastically with the pendulum. The bullet has a mass of 5 grams and a speed of 100 m/s. The pendulum has a mass of 2 kg and a length of 0.5 m. To what angle (from vertical) does the pendulum swing?
Answer #1

0.500 0.5 Cose h=0.5-0.5 cose , h=0.5(1-Coso) briven that m=5gm 5x100 v=toomis - M - 2kg Since cotho it is Ballistic Pendulum